A spherical balloon is deflated so that its volume is decreasing at a rate of 3 ftଷ/min. How fast is the diameter of the balloon decreasing when the radius is 2 ft? 


1
Expert's answer
2021-09-27T16:33:48-0400
"V=\dfrac{4}{3}\pi R^3=\dfrac{1}{6}\pi D^3"

Differentiate both sides with respect to "t" and use the Chain Rule


"\dfrac{dV}{dt}=\dfrac{1}{6}\pi(3D^2)\dfrac{dD}{dt}"

"\dfrac{dD}{dt}=\dfrac{2}{\pi D^2}\dfrac{dV}{dt}"

Given "\dfrac{dV}{dt}=-3\ {ft}^3/min"

When "R=2\ ft, D=4\ ft"


"\dfrac{dD}{dt}=\dfrac{2}{\pi (4\ ft)^2}(-3\ ft^3/min)=-\dfrac{3}{8\pi }\ ft/min"

The diameter of the balloon is decreasing at a rate of "\dfrac{3}{8\pi }\approx0.12" ft/min when the radius is 2 ft.